A student got 150 marks out of 200 in maths and got 120 marks out of 180 in science. In which subject did the student perform better?

Grade:12

2 Answers

Aditya Gupta
3 years ago
percentage marks in maths: (150/200)*100= 75%
percentage marks in science: (120/180)*100= 66.666%
so, he performed better in MATHS.
KINDLY APPROVE :))
 
Harshit Singh
3 years ago
Dear Student
Expressing the marks in the form of ratios.

For maths, ratio = 150/200 = ¾

For science, ratio = 120/180= ⅔

Here, the ratio ¾ shows 75% (¾ × 100 = 75) and the ratio ⅔ shows 66.6% (⅔ × 100 = 66.6).

Thus, the student performed better in the maths exam.


Thanks